What is the Jacobian matrix J(r, θ) for the polar coordinate transformation, given that x=rcos⁡θ and y=rsin⁡θ.



Select one:



A. cos⁡θ-rsin⁡θ-sin⁡θrcos⁡θ




B. -rcos⁡θsin⁡θsin⁡θ-rcos⁡θ




C. cos⁡θ-rsin⁡θsin⁡θ-rcos⁡θ




D. cos⁡θrsin⁡θsin⁡θrcos⁡θ


1
Expert's answer
2022-04-06T09:31:57-0400

"J=\\left[ \\begin{matrix}\t\\frac{\\partial}{\\partial r}\\left( r\\cos \\theta \\right)&\t\t\\frac{\\partial}{\\partial \\theta}\\left( r\\cos \\theta \\right)\\\\\t\\frac{\\partial}{\\partial r}\\left( r\\sin \\theta \\right)&\t\t\\frac{\\partial}{\\partial \\theta}\\left( r\\sin \\theta \\right)\\\\\\end{matrix} \\right] =\\left[ \\begin{matrix}\t\\cos \\theta&\t\t-r\\sin \\theta\\\\\t\\sin \\theta&\t\tr\\cos \\theta\\\\\\end{matrix} \\right] \\\\"

Only one '-' sign, no correct variant